Given that matrix A has 1has the eigenvalue with the corresponding eigenvector (1,-3) and 5 as the second eigenvalue with (1,1) as the corresponding eigenvector. (I) find the matrix A (Ii) D the diagonal matrix such that P^-1 AP=D


1
Expert's answer
2021-07-12T16:08:01-0400

Let

"A=\begin{pmatrix}
 a & b \\
 c & d
\end{pmatrix}"


"\begin{vmatrix}
 a-\lambda & b \\
 c & d-\lambda
\end{vmatrix}=0"

"(a-\lambda)(d-\lambda)-bc=0"

"\lambda^2-(a+d)\lambda+ad-bc=0"


"\lambda_1=1, v_1=\begin{pmatrix}
 1 \\
 -3
\end{pmatrix}"

"\lambda_2=5, v_1=\begin{pmatrix}
 1 \\
 1
\end{pmatrix}"

By Vieta Theorem


"a+d=1+5"

"ad-bc=1(5)"

"Av_i=\lambda_iv_i, i=1,2"

"\begin{pmatrix}
 a-1 & b \\
 c & d-1
\end{pmatrix}\begin{pmatrix}
 1 \\
 -3
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0
\end{pmatrix}"

"a-1-3b=0""c-3d+3=0"

"a=3b+1""c=3d-3"


"\begin{pmatrix}
 a-5 & b \\
 c & d-5
\end{pmatrix}\begin{pmatrix}
 1 \\
 1
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0
\end{pmatrix}"

"a-5+b=0""c+d-5=0"

"a=-b+5""c=-d+5"

"3b+1=-b+5=>b=1, a=4"

"3d-3=-d+5=>d=2, c=3"

"A=\begin{pmatrix}
 4 & 1 \\
 3 & 2
\end{pmatrix}"

"P=\begin{pmatrix}
 1 & 1 \\
 -3 & 1
\end{pmatrix}, P^{-1}=\begin{pmatrix}
 0.25 & -0.25 \\
 0.75 & 0.25
\end{pmatrix}"

"D=P^{-1}AP"

"D=\begin{pmatrix}
 1 & 0 \\
 0 & 5
\end{pmatrix}"